# Correcting the LogQ Correction: Revisiting Sampled Softmax for Large-Scale Retrieval

### v1

@@ 1. Introduction
In this paper, we identify and address this issue with a new formulation of logQ correction.
@@ 3.3. Correcting the Correction
In the standard logQ correction, positive item $p$ is used together with sampled items to produce Monte-Carlo estimate of the expectation over the full catalog. However, in contrast to negatives, it is not sampled from the proposal distribution $Q(d)$ — for a given training instance, it is deterministically present with probability 1. This mismatch introduces additional bias.
@@
The introduced weight $w_{up}$ reflects the model’s confidence: if the model is confident that the positive item $p$ is the correct one, the weight will be small.
To estimate $P_{\theta}(p\mid u)$ needed for the weight, importance sampling is applied once more:
@@ 4. Experiments
We evaluate our method in both academic and industrial setups to demonstrate its robustness across different environments. The industrial setup is based on a large-scale production system, reflecting realistic constraints and item distributions.
@@ 4.1. Academic Setup
    • Mixed negative sampling(Yang et al., 2020), which combines both sources above.
@@
For gBCE and sampled softmax, we sample 256 negatives. In the mixed negative sampling (MNS)(Yang et al., 2020) setup, 128 negatives are sampled uniformly and 128 from in-batch items. Importantly, for mixed sampling, $\log Q$ values are based on unigram frequencies and not adjusted for the uniform portion.
@@ 4.2. Industrial Setup
For evaluation, we use user interactions from the day following the training period, measuring performance separately for consumption and engagement signals. The retrieval index includes the top 2 million most popular items from training, along with all positive user-item interactions from the test day. User embeddings are computed from data preceding the test day. We report Recall@k for $k\in\{10,100,1000\}$. The goal is to retrieve, for each user, as many of their next-day positive interactions as possible within a relatively small candidate set.
@@
In this industrial setting, the model serves as a retrieval model, whose primary goal is to produce a large pool of relevant items (typically thousands) for downstream reranking. Consequently, recall is the key metric: improvements in Recall@100 or Recall@1000 translate directly into measurable gains in downstream system performance. In contrast, metrics such as NDCG, while informative for final ranking tasks, are less meaningful in the retrieval stage and can be misleading when a separate model handles ranking. This practice is common across large-scale retrieval systems in industry (Yi et al., 2019; Yang et al., 2020; Li et al., 2021; Pancha et al., 2022; Baltescu et al., 2022; Nigam et al., 2019).
@@ 4.3. Results
Table 1. Recall@20 and NDCG@20 performance for different models comparing leave-one-out and time split validation on the public datasets
@@
Table 1 shows that no single loss dominates across all datasets and evaluation setups. Full softmax performs best only on Steam (leave-one-out), while uniform sampled softmax unexpectedly leads on Steam (temporal split). On MovieLens, logQ-based corrections achieve the strongest performance under temporal-split evaluation, and are competitive under leave-one-out as well — falling slightly behind gBCE in Recall@20 but outperforming it in NDCG@20, with minimal margin.
In-batch negatives without any correction result in substantially degraded metrics. Applying logQ correction recovers most of the gap, and combining it with MNS (mixed negative sampling) further improves performance — often pushing logQ-corrected models to the top. Our improved logQ correction with MNS yields consistent gains over the standard variant: the improvement is large on MovieLens (leave-one-out) and across all metrics on Steam, while slightly trailing standard logQ only on MovieLens (temporal split).
@@ Appendix A Investigating Standard LogQ Correction
$\mathcal{L}_{\text{softmax}}(u,p)=-\log\frac{e^{f_{\theta}(u,p)}}{\sum_{d\in\mathcal{D}}e^{f_{\theta}(u,d)}}.$
@@
We next employ the importance sampling to replace $P_{\theta}(d\mid u)$ in the expectation with the proposal distribution $Q(d)$, followed by Monte-Carlo approximation of the expectation:
@@
where $d_{i}\sim Q(d)$. Subsequently, we approximate the denominator term using the same importance sampling technique combined with Monte-Carlo estimation:
@@
A simpler way to derive the same estimate is to use weighted importance sampling instead of applying standard one twice:
@@ Appendix B Correcting the Correction: Full Derivation
We next employ weighted importance sampling to replace $P_{\theta}(d\mid u,d\neq p)$ distribution with the proposal $Q^{\prime}(d)$ over $\mathcal{D}\setminus\{p\}$ which is different from $Q(d)$ in that it explicitly excludes the positive item $p$ from being sampled.
@@
where sg denotes the stop-gradient operation (no gradient flows through $P_{\theta}(p\mid u)$).
To estimate $P_{\theta}(p\mid u)$, importance sampling is applied once more:

### v2

@@ 1. Introduction
We address this inconsistency by introducing a corrected formulation of logQ correction.
@@ 3.3. Correcting the Correction
In the standard logQ correction, positive item $p$ is used together with sampled items to produce Monte Carlo estimate of the expectation over the full catalog. However, in contrast to negatives, it is not sampled from the proposal distribution $Q(d)$ — for a given training instance, it is deterministically present with probability 1. This mismatch introduces additional bias.
@@
The introduced weight $w_{up}$ reflects the model’s confidence in distinguishing the positive item $p$ from the negatives: if the model assigns a high score to $p$ relative to the negatives, the weight will be small.
To estimate $P_{\theta}(p\mid u)$ for $w_{up}$, importance sampling is applied once more:
@@ 4. Experiments
We evaluate our method in both academic and industrial setups to demonstrate its robustness across different environments. The industrial setup is based on a large-scale production system.
@@ 4.1. Academic Setup
    • Mixed negative sampling (MNS)(Yang et al., 2020), which combines both sources above.
@@
For gBCE and sampled softmax, we sample 256 negatives. In the MNS setup, 128 negatives are sampled uniformly and 128 from in-batch items. Importantly, for mixed sampling, $\log Q$ values are based on unigram frequencies and not adjusted for the uniform portion.
@@ 4.2. Industrial Setup
For evaluation, we use user-item interactions from the day following the training period, measuring performance separately for consumption and engagement signals. The retrieval index includes the top 2 million most popular items from training, along with all positive user-item interactions from the test day. User embeddings are computed from data preceding the test day. We report Recall@k for $k\in\{10,100,1000\}$. The goal is to retrieve, for each user, as many of their next-day positive interactions as possible within a relatively small candidate set.
@@
In our industrial setup, the model serves as a retrieval model, whose primary goal is to produce a large pool of relevant items (typically thousands) for downstream reranking. Consequently, recall is the key metric: improvements in Recall@100 or Recall@1000 translate directly into measurable gains in downstream system performance. In contrast, metrics such as NDCG, while informative for final ranking tasks, are less meaningful in the retrieval stage and can be misleading when a separate model handles ranking. This practice is common across large-scale retrieval systems in industry (Yi et al., 2019; Yang et al., 2020; Li et al., 2021; Pancha et al., 2022; Baltescu et al., 2022; Nigam et al., 2019).
@@ 4.3. Results
Table 1. Recall@20 and NDCG@20 performance for different models comparing leave-one-out and temporal split validation on the public datasets
@@
Table 1 shows that no single loss dominates across all datasets and evaluation setups. Full softmax performs best only on Steam (leave-one-out), while uniform sampled softmax unexpectedly leads on Steam (temporal split). On MovieLens, logQ corrections achieve the strongest performance under temporal-split evaluation, and are competitive under leave-one-out as well — falling slightly behind gBCE in Recall@20 but outperforming it in NDCG@20, with minimal margin.
In-batch negatives without any correction result in substantially degraded metrics. Applying logQ correction recovers most of the gap, and combining it with MNS further improves performance — often pushing logQ-corrected models to the top. Our improved logQ correction with MNS yields consistent gains over the standard variant: the improvement is large on MovieLens (leave-one-out) and across all metrics on Steam, while slightly trailing standard logQ only on MovieLens (temporal split).
@@ Appendix A Investigating Standard LogQ Correction
$\mathcal{L}_{\text{softmax}}(u,p)=-\log\frac{e^{f_{\theta}(u,p)}}{\sum_{d\in\mathcal{D}}e^{f_{\theta}(u,d)}},$
@@
We next employ the importance sampling to replace $P_{\theta}(d\mid u)$ in the expectation with the proposal distribution $Q(d)$:
@@
where $d_{i}\sim Q(d)$. Subsequently, we approximate the denominator term using the same importance sampling technique:
@@
Weighted importance sampling.
A simpler way to derive the same estimate is to use weighted importance sampling:
@@ Appendix B Correcting the Correction: Full Derivation
We next employ weighted importance sampling to replace $P_{\theta}(d\mid u,d\neq p)$ distribution with the proposal $Q^{\prime}(d)$ over $\mathcal{D}\setminus\{p\}$ which is different from $Q(d)$ in that it explicitly excludes the positive item $p$ from being sampled:
@@
where sg denotes the stop-gradient operation (no gradient flows through $P_{\theta}(p\mid u)$). To estimate $P_{\theta}(p\mid u)$, importance sampling is applied once more: